Calculate the pH and pOH of the following solution and identify each solution as acidic, basic or neutral. (with solutions)


1. [H3O+] = 1.0 x 10-9 M

2. [OH-] = 1.0 x 10-8 M

3. [H3O+] = 0.0010 M

4. [OH-] = 0.010 M




1
Expert's answer
2022-03-23T14:51:02-0400

In the question molarity of different solution is given

and we know that

"pH= -log[H_3O^+]"


"pOH= -log[OH^{-1}]"


1 . "pH= -log[1.0\times 10^{-9}]"


pH= 9


Since, its pH value is 9 so it is mild basic


2 . "pOH= -log[1.0\times 10^{-8}]=8"


pH = 14-8= 6

so it is weak acid


3 . "pH= -log[1.0\times 10^{-3}]=3"


It is acidic


4 . "pOH= -log[1.0\times 10^{-2}]= 2"


pH = 12 it is basic